Solve the system of equations and give the value of x.
8x = 23 + 7y
2x + 4y = 0

Found 2 solutions by tommyt3rd, lenny460:
Answer by tommyt3rd(5050)

2x+4y=0
x=-2y


8x=23+7y
8(-2y)=23+7y
-16y-7y=23
y=-1

x=-2
solution: (-2,-1)

Answer by lenny460(1073)

Solve the system of equations and give the value of x.
8x = 23 + 7y
2x + 4y = 0

2x + 4y = 0
Divide each item by 2
Therefore:
x + 2y = 0
Subtract 2y from each side of the equation.
x + 2y - 2y = 0 - 2y
x = -2y

Rearrange equation 1:
8x = 23 + 7y
Subtract 7y from each side of the equation
8x - 7y = 23 + 7y - 7y
8x - 7y = 23
Substitute x = -2y into equation 1
8(-2y) - 7y = 23
-16y - 7y = 23
-23y = 23
Divide each side by -23
y = -1

We already have:
x = -2y
Therefore:
x = -2(-1)
x = 2

Check:
8x - 7y = 23
8(2) - 7(-1) = 23
16 - (-7) = 23
- (-7) = +7
16 + 7 = 23
23 = 23


Check:
2x + 4y = 0
2(2) + 4(-1) = 0
4 + (-4) = 0
4 - 4 = 0




Answer:
x = 2 and
y = -1








Lennox Obuong
Algebra Tutor
Nairobi, Kenya
Email: lennoxobuong@yahoo.com